The sum of James and tom 2 years ago Was 40. In two years time from now, the ages of James will be 3times than of john. Find their ages now

Answer by Boreal(15235)

ages now are x for James and y for Tom
(x-2)+(y-2)=40, x+y-4=40, x+y=44
3(y+2)=x+2, James will be 3 times that of Tom in 2 years.
---------------
x+y=44, rewriting.
x+2=3y+6, x-3y-4=0
3x+3y=132
x-3y=4
4x=136
x=34. James is 34 now. Two years ago, he was 32 and Tom was 8. Therefore, Tom is 10 now.
In two years, James will be 36 and Tom 12, a 3:1 ratio.